Let f be a smooth function. Calculate the curvature and the torsion of the curve that is the intersection of x = y and z = f(x).


1
Expert's answer
2022-06-27T05:01:51-0400

Solution:

Formula of calculation curvature (for x=y):

"k=\frac{x'z''-z'x''}{(x'^{2}+z'^{2})};" here "z= f(x); \space and \space z'=\frac{df(x)}{dx}; \space z''=\frac{d^{2}f(x)}{dx^{2}};"

so, "x'=1 \space and \space x''=0;" "k=\frac{z''}{(1+z'^{2})};"

For torsion:

"\tau=\frac{x'''(y'z''-y''z')+y'''(x''z'-x'z'')+z'''(x'y''-x''y')}{(y'z''-y''z')^{2}+(x''z'-x'z'')^{2}+(x'y''-x''y')^{2}};"

"x''=0\space and\space y=0;"

So, "\tau=0."

Answer:

"k=\frac{z''}{(1+z'^{2})};"

"\tau=0."